# Freeze-in and Freeze-out production of Higgs Portal Majorana Fermionic Dark Matter during and after Reheating

### v1

@@ III.1.2 Freeze-out
During reheating, inflaton decays into SM particles; leading to non-conservation of entropy. This physical situation give rise to two distinct scenarios:
(a) Freeze out after reheating ($T_{\rm fo}<T_{\rm re}$): This is known as the standard freeze-out scenario, where DM freeze-out occurs well after reheating, in the standard radiation-dominated era. After the freeze-out, the comoving number density $N_{\chi}=n_{\chi}A^{3}$ will be much larger than the comoving equilibrium number density $N_{\rm eq}^{r}$.Therefore, once freeze-out occurs, $N_{\rm eq}$ can be neglected in comparison to $N_{\rm{\chi}}$ and from the Eq. 24,
${}\frac{N_{\chi}(A)}{dA}=-\frac{\langle\sigma v\rangle}{H}A^{-4}\,N^{2}_{\chi}.$ (38)
Taking into account the radiation-dominated era and integrating from freeze-out point ($T_{\rm fo}$) to the present time ($T_{0}$),we get the present-day comoving number density ($N_{0}$)
$N_{0}=\frac{16\,\pi\,\Lambda^{2}}{3}\,\frac{m_{\chi}\,H_{\rm re}}{T_{\rm re}}A_{\rm re}^{3}\left(\frac{T_{\rm re}}{T_{\rm fo}}\right)^{2}\,.$ (39)
The freeze-out temperature $T_{\rm fo}$ can be written as (from Eq. 37),
$T_{\rm fo}=2\,m_{\rm\chi}\frac{1}{W_{-1}\left[2\,m_{\rm\chi}\,\mathcal{K}^{-2}_{0}\right]}~{}~{}~{}\mbox{where}~{}~{}~{}\mathcal{K}_{0}=\frac{16\,\pi\,\Lambda^{2}}{3\sqrt{2}}\left(\frac{\epsilon\,\pi}{m_{\rm\chi}}\right)^{1/2}\,,$ (40)
where $W_{-1}$ is the Lambert function with branch $-1$. Therefore, the current abundance can be written as,
$\Omega_{\chi}h^{2}=\Omega_{\rm R}h^{2}\frac{m_{\rm\chi}}{\epsilon\,T_{0}}\frac{N_{0}}{T_{\rm re}^{3}\,A_{\rm re}^{3}}=\Omega_{\rm R}h^{2}\frac{16\,\pi}{3\,\sqrt{3\,\epsilon}}\frac{\Lambda^{2}}{T_{0}\,m_{p}}\left(\frac{m_{\chi}}{T_{\rm fo}}\right)^{2}\,.$ (41)
Interestingly, both the freeze-out temperature and relic abundance are intriguingly independent of the reheating temperature $T_{\rm re}$, a remarkable contrast to the freeze-in mechanism where $T_{\rm re}$ plays a pivotal role.
Figure 5: The evolution of the comoving DM number density for freeze-out as a function of the scale factor $A$ (bottom x-axis) and the bath temperature $T$ (top x-axis) for $w_{\rm\phi}=0.0$ (left), $w_{\rm\phi}=0.5$ (right) with $T_{\rm re}=10^{9}$ GeV. Each plot presents two different cases: freeze-out after reheating $T_{\rm fo}<T_{\rm re}$ (solid), and freeze-out during reheating $T_{\rm fo}>T_{\rm re}$ (dashed). The vertical red line corresponds to $T\sim T_{\rm fo}$, and the vertical black line corresponds to $T=T_{\rm re}$.
@@
$T_{\rm fo}=\frac{2\,m_{\rm\chi}}{5-2\,j}\frac{1}{W_{-1}\left[\frac{2\,m_{\rm\chi}}{5-2\,j}\,\mathcal{K}^{\frac{2}{2\,j-5}}\right]}~{}~{}~{}~{}\mbox{where}~{}~{}~{}\mathcal{K}=\frac{8\,\pi(2\,\pi)^{3/2}}{3\,g}\frac{\Lambda^{2}\,H_{\rm re}}{\sqrt{m_{\rm\chi}}}T_{\rm re}^{-j}\,,j=\frac{4\,(1+w_{\rm\phi})}{1+3\,w_{\rm\phi}}\,.$ (44)
@@
(a) Freeze-out after reheating ($T_{\rm fo}<T_{\rm re}$): In this case, dark matter freeze-out takes place well after reheating, provided that the DM mass satisfies $m_{\rm\chi}<25\,T_{\rm re}$. The resulting relic abundance, $\Omega_{\chi}h^{2}$, is independent of the inflaton equation of state ($w_{\rm\phi}$), reheating temperature ($T_{\rm re}$), and DM mass ($m_{\rm\chi}$). The observed relic abundance is typically achieved for a coupling scale of $\Lambda\simeq\mathcal{O}(10^{3})$ GeV, irrespective of the DM mass (solid line).
(b) Freeze-out during reheating ($T_{\rm fo}>T_{\rm re}$): If the DM mass is $m_{\rm\chi}\gtrsim 25\,T_{\rm re}$, freeze-out occurs during the reheating phase. In this case, the relic abundance scales as $\Omega_{\chi}h^{2}\propto\Lambda^{2}\,m_{\rm\chi}^{(7w_{\rm\phi}-3)/(1+3w_{\rm\phi})}$ (from Eq. 43), assuming $T_{\rm fo}\propto m_{\rm\chi}$. Consequently, the slope of the $\Lambda$ vs. $m_{\rm\chi}$ curve varies depending on the inflaton EoS $w_{\rm\phi}$; for $w_{\rm\phi}=0.0$, the coupling scale follows $\Lambda\propto m_{\rm\chi}^{3/2}$, meaning that $\Lambda$ increases with $m_{\rm\chi}$. For $w_{\rm\phi}=1/3$, the relation becomes $\Lambda\propto m_{\rm\chi}^{1/6}$, showing a milder increase. However, for $w_{\rm\phi}=0.5$, the trend reverses, with $\Lambda$ decreasing as $m_{\rm\chi}$ increases, following $\Lambda\propto m_{\rm\chi}^{-0.1}$. All these behaviors are illustrated with red dashed lines in Fig. 6.
@@ III.2 DM production after EWSB
${}\dot{n}_{\rm x}+3Hn_{\rm x}+\langle\sigma v\rangle(n_{\rm x}^{2}-n_{\rm eq}^{2})+\langle\Gamma_{h}\rangle n_{\rm h}\left(1-\frac{n^{2}_{\rm x}}{n^{2}_{\rm eq}}\right)=0\,,$ (45)
@@
Figure 7: The evolution of the co-moving DM number density for freeze-in mechanism as a function of the scale factor $A$ (bottom x-axis) and the bath temperature $T$ (top x-axis) for $w_{\rm\phi}=0.0$. The vertical red lines indicate $T\sim T_{\rm fo}$, while the vertical solid and dashed black lines represent $T=T_{\rm re}$ and $T=T_{\rm EW}$, respectively.
@@ III.2.1 Freeze-in
We have analyzed the behavior of the DM parameter space $(\Lambda,m_{\rm\chi})$ in Fig. 8 (top panel) for the freeze-in mechanism in the post-EWSB scenario. We found that if $w_{\rm\phi}<3/7$ and reheating temperature $T_{\rm re}<T_{\rm EW}$, EWSB can affect freeze-in production only if the DM mass below $\mathcal{O}(500)$ GeV. In Fig. 8, we consider $T_{\rm re}=10$ GeV ($<T_{\rm EW}$) for two different EoS $w_{\rm\phi}=(0,1/3)$. A striking qualitative difference emerges between DM masses above and below $m_{\rm h}/2$, with the most noticeable feature being the sharp drop at the Higgs resonance $m_{\rm\chi}\approx m_{\rm h}/2$. For $m_{\rm\chi}<m_{\rm h}/2$, DM is dominantly produced via Higgs decays, leading to an expected relic abundance proportional to $m_{\rm\chi}$, as freeze-in occurs around $T\sim m_{\rm h}/2$ and the slope is $\Lambda\propto m_{\rm\chi}^{1/2}$ (see solid line). In contrast, for $m_{\rm\chi}>m_{\rm h}/2$, production is dominated by 2-to-2 scatterings, which changes the slope of the $\Lambda$ vs. $m_{\rm\chi}$ relation (dashed line). Due to the complexity of the interactions , an analytical estimation of the slope in this regime is not possible. Note that when $w_{\rm\phi}>3/7$, all DM masses are produced before EWSB irrespective of $T_{\rm re}$, and its behavior is the same as the pre-EWSB case, therefore we do not show the parameter space for $w_{\rm\phi}=0.5$.
@@ III.2.2 Freeze-out
We have illustrated the behavior of the DM parameter space $(\Lambda,m_{\rm\chi})$ in Fig. 8 for the freeze-out mechanism in the post-EWSB scenario. This parameter space is obtained by numerically solving Eq. 45 together with Eq. 5. We find that any DM mass $m_{\rm\chi}\gtrsim\mathcal{O}(4)$ TeV freezes out before EWSB; otherwise, freeze-out occurs after EWSB. Thus, EWSB scenario can significantly influence the DM freeze-out, if the DM mass is below $\mathcal{O}(4)$ TeV, for any $w_{\rm\phi}$ and $T_{\rm re}$. The impact of reheating on freeze-out production is shown with red dashed lines. The solid red line corresponds to the standard freeze-out scenario, in which DM freezes out in the radiation-dominated epoch.
@@ IV DM parameter space and its constraints
In the previous section, we provided a detailed analysis of DM production and the behavior of its parameter space. In this section, we extend our discussion to explore the complete DM parameter space along with its relevant constraints. Figure. 9 illustrates the complete DM parameter space, where we consider six distinct reheating temperatures: $T_{\rm re}=(T_{\rm BBN},10,10^{5},10^{9},10^{12},10^{15})$ GeV, covering the entire range from the minimum to the maximum reheating temperatures. The colored shaded regions represent different constraints. The red-shaded region corresponds to the bound on $T_{\rm re}$. The light blue region represents the parameter space where the Higgs-Majorana interaction is Planck-suppressed ($\Lambda>M_{\rm pl}$), supported by the WIMPZILLA hypothesis Kolb and Long (2017). Since $T_{\rm max}$ is the maximum temperature during reheating, the effective theory of DM described in Eq. 21 is valid only if $\Lambda\geq T_{\rm max}$ Garcia et al. (2020); Barman et al. (2024a); the brown-shaded region violates this condition. This constraint rules out the entire freeze-out parameter space (see right plot), while in the freeze-in scenario, only certain regions of the DM parameter space are excluded. Additionally, DM masses below $\mathcal{O}(10^{-5})$ GeV are excluded by Lyman-$\alpha$ Iršič et al. (2017); Ballesteros et al. (2021) constraints for freeze-in (see Appendix-A). The small black circle represents the coincidence point of freeze-in and freeze-out mechanisms. The corresponding mass at this point is the maximum DM mass capable of producing the correct relic abundance. For masses above this value, the DM becomes underabundant today.
@@
Figure 10: Comparison of the freeze-in DM parameter space $(\Lambda^{-1},m_{\rm\chi})$ with the sensitivity of current and future detection experiments. The three red lines correspond to different reheating temperatures: $T_{\rm re}=(T_{\rm BBN},0.5,10)$ GeV. Shaded regions indicate experimental sensitivities: XENONnT (green), LZ (purple), XLZD (orange), and LHC constraints (magenta), while the brown shaded area represents excluded parameter space ($\Lambda<T_{\rm max}$). The blue dashed line (“$\nu$ fog”) marks the neutrino background limit for direct DM detection.”

### v2

@@ III.1.2 Freeze-out
During reheating, the inflaton decays into SM particles, leading to non-conservation of entropy. This physical situation gives rise to two distinct scenarios:
(a) Freeze out after reheating ($T_{\rm fo}<T_{\rm re}$): This is known as the standard freeze-out scenario, where DM freeze-out occurs well after reheating, in the standard radiation-dominated era. After the freeze-out, the comoving number density $N_{\chi}=n_{\chi}A^{3}$ will be much larger than the comoving equilibrium number density $N_{\rm eq}^{r}$. Therefore, once freeze-out occurs, $N_{\rm eq}$ can be neglected [^3] Our main results are based on a complete numerical solution of the Boltzmann equation Eq. 24, while the analytical solution of Eq. 38 is provided to give intuitive, physical insight into the scaling behavior of DM abundance. in comparison to $N_{\rm{\chi}}$ and from the Eq. 24,
$\frac{dN_{\chi}(A)}{dA}=-\frac{\langle\sigma v\rangle}{H}A^{-4}\,N^{2}_{\chi}.$ (38)
Taking into account the radiation-dominated era and integrating from the freeze-out point ($T_{\rm fo}$) to the present time ($T_{0}$), we get the present-day comoving number density ($N_{0}$)
$N_{0}=\frac{16\,\pi\,\Lambda^{2}}{3}\,\frac{m_{\chi}\,H_{\rm re}}{T_{\rm re}}A_{\rm re}^{3}\left(\frac{T_{\rm re}}{T_{\rm fo}}\right)^{2}=\frac{16\,\pi\,\sqrt{\epsilon}\,\Lambda^{2}}{3\,\sqrt{3}}\,\frac{m_{\rm\chi}\,T_{\rm fo}}{M_{\rm pl}}\,A^{3}_{\rm fo}\,.$ (39)
We replaced $H_{\rm re}$ in terms of $T_{\rm re}$ (using Eq. 20), to obtain the final expression. The freeze-out temperature $T_{\rm fo}$ can be written as (from Eq. 37),
$T_{\rm fo}=2\,m_{\rm\chi}\,\frac{1}{W_{-1}\left[2\,m_{\rm\chi}\,\mathcal{K}^{-2}_{0}\right]}~~~\mbox{where}~~~\mathcal{K}_{0}=\frac{16\,\pi^{2}\,\Lambda^{2}}{3\,g_{\chi}\,M_{\rm pl}}\left(\frac{2\,\epsilon\,\pi}{3\,m_{\rm\chi}}\right)^{1/2}\,$, (40)
where $W_{-1}$ is the Lambert function with branch 1-1. Therefore, the current abundance can be written as,
$\Omega_{\chi}h^{2}=\Omega_{\rm R}h^{2}\frac{m_{\rm\chi}}{\epsilon\,T_{0}}\frac{N_{0}}{T_{\rm re}^{3}\,A_{\rm re}^{3}}=\Omega_{\rm R}h^{2}\frac{16\,\pi}{3\,\sqrt{3\,\epsilon}}\frac{\Lambda^{2}}{T_{0}\,M_{\rm pl}}\left(\frac{m_{\chi}}{T_{\rm fo}}\right)^{2}\,.$ (41)
Interestingly, both the freeze-out temperature and relic abundance are intriguingly independent of the reheating temperature $T_{\rm re}$, a remarkable contrast to the freeze-in mechanism where $T_{\rm re}$ plays a pivotal role. This confirms that the standard freeze-out scenario is independent of the dynamics of reheating.
Figure 5: The evolution of the comoving DM number density for freeze-out as a function of the scale factor $A$ (bottom x-axis) and the bath temperature $T$ (top x-axis) for $w_{\rm\phi}=0.0$ (left), $w_{\rm\phi}=0.5$ (right) with $T_{\rm re}=10^{9}$ GeV. Each plot presents two different cases: freeze-out after reheating $T_{\rm fo}<T_{\rm re}$ (solid line), and freeze-out during reheating $T_{\rm fo}>T_{\rm re}$ (dashed line). In both plots, $\Lambda=10^{8}$ GeV. The vertical red line corresponds to $T\sim T_{\rm fo}$, and the vertical black line corresponds to $T=T_{\rm re}$.
@@
$T_{\rm fo}=\frac{2\,m_{\rm\chi}}{5-2\,j}\frac{1}{W_{-1}\left[\frac{2\,m_{\rm\chi}}{5-2\,j}\,\mathcal{K}^{\frac{2}{2\,j-5}}\right]}~~~~\mbox{where}~~~\mathcal{K}=\frac{8\,\pi(2\,\pi)^{3/2}}{3\,g_{\rm\chi}}\frac{\Lambda^{2}\,H_{\rm re}}{\sqrt{m_{\rm\chi}}}T_{\rm re}^{-j}\,,j=\frac{4\,(1+w_{\rm\phi})}{1+3\,w_{\rm\phi}}\,.$ (44)
@@
(a) Freeze-out after reheating ($T_{\rm fo}<T_{\rm re}$): In this case, dark matter freeze-out takes place well after reheating, provided that the DM mass satisfies $m_{\rm\chi}<25\,T_{\rm re}$ Bernal and Xu (2022); Silva-Malpartida et al. (2023). This condition follows from the fact that freeze-out typically occurs at a temperature $T_{\rm fo}\simeq m_{\chi}/x_{f}$, with $x_{f}\sim 20$-25. The resulting relic abundance, $\Omega_{\chi}h^{2}$, is independent of the inflaton equation of state ($w_{\rm\phi}$), reheating temperature ($T_{\rm re}$), and DM mass ($m_{\rm\chi}$). Although $m_{\chi}$ appears explicitly in the expression of $T_{\rm fo}$ (Eq. 40), the Lambert function ensures that $T_{\rm fo}$ scales approximately linearly with $m_{\chi}$. Consequently, in expressions such as Eq. (33), where $(m_{\chi}/T_{\rm fo})^{2}$ appears, the explicit dependence on $m_{\chi}$ effectively cancels. This explains why the relic abundance $\Omega_{\chi}h^{2}$ is approximately independent of the DM mass 444with a small logarithmic dependence on the DM mass, despite the apparent appearance of $m_{\chi}$ in the formula. The observed relic abundance is typically achieved for a coupling scale of $\Lambda\simeq\mathcal{O}(10^{3})$ GeV, irrespective of the DM mass (solid line).
(b) Freeze-out during reheating ($T_{\rm fo}>T_{\rm re}$): If the DM mass is $m_{\rm\chi}\gtrsim 25\,T_{\rm re}$ Bernal and Xu (2022); Silva-Malpartida et al. (2023), freeze-out occurs during the reheating phase. In this case, the relic abundance scales as $\Omega_{\chi}h^{2}\propto\Lambda^{2}\,m_{\rm\chi}^{(7w_{\rm\phi}-3)/(1+3w_{\rm\phi})}$ (from Eq. 43), assuming $T_{\rm fo}\propto m_{\rm\chi}$. Consequently, the slope of the $\Lambda$ vs. $m_{\rm\chi}$ curve varies depending on the inflaton EoS $w_{\rm\phi}$; for $w_{\rm\phi}=0.0$, the coupling scale follows $\Lambda\propto m_{\rm\chi}^{3/2}$, meaning that $\Lambda$ increases with $m_{\rm\chi}$. For $w_{\rm\phi}=1/3$, the relation becomes $\Lambda\propto m_{\rm\chi}^{1/6}$, showing a milder increase. However, for $w_{\rm\phi}=0.5$, the trend reverses, with $\Lambda$ decreasing as $m_{\rm\chi}$ increases, following $\Lambda\propto m_{\rm\chi}^{-0.1}$. All these behaviors are illustrated with red dashed lines in Fig. 6.
We note that some of the dark matter (DM) masses shown in Fig. 6, presented for illustrative purposes, exceed the unitarity bound for thermal WIMPs, $m_{\chi}\gtrsim 100~\mathrm{TeV}$. These examples are included solely to illustrate the behavior of the DM parameter space during freeze-out after reheating. Nevertheless, the qualitative features of the $\Lambda$-$m_{\chi}$ relation remain valid even beyond the unitarity limit.
@@ III.2 DM production after EWSB
$\dot{n}_{\chi}+3Hn_{\chi}+\langle\sigma v\rangle(n_{\chi}^{2}-n_{\rm eq}^{2})+\langle\Gamma_{h}\rangle n_{\rm h}\left(1-\frac{n^{2}_{\chi}}{n^{2}_{\rm eq}}\right)=0\,$, (45)
@@
Figure 7: The evolution of the co-moving DM number density for freeze-in mechanism as a function of the scale factor $A$ (bottom x-axis) and the bath temperature $T$ (top x-axis) for $w_{\rm\phi}=0.0$. The vertical red lines indicate $T\sim T_{\rm fi}$, while the vertical solid and dashed black lines represent $T=T_{\rm re}$ and $T=T_{\rm EW}$, respectively.
@@ III.2.1 Freeze-in
We have analyzed the behavior of the DM parameter space $(\Lambda,m_{\rm\chi})$ in Fig. 8 (top panel) for the freeze-in mechanism in the post-EWSB scenario. We found that if $w_{\rm\phi}<3/7$ and reheating temperature $T_{\rm re}<T_{\rm EW}$, EWSB can affect freeze-in production only if the DM mass below $\mathcal{O}(500)$ GeV. In Fig. 8, we consider $T_{\rm re}=10$ GeV ($T_{\rm EW}$) for two different EoS $w_{\rm\phi}=(0,1/3)$. A striking qualitative difference emerges between DM masses above and below $m_{\rm h}/2$, with the most noticeable feature being the sharp drop at the Higgs resonance $m_{\rm\chi}\approx m_{\rm h}/2$. For $m_{\rm\chi}<m_{\rm h}/2$, DM is dominantly produced via Higgs decays, leading to an expected relic abundance proportional to $m_{\rm\chi}$, as freeze-in occurs around $T\sim m_{\rm h}/2$ and the slope is $\Lambda\propto m_{\rm\chi}^{1/2}$ (see solid line). In contrast, for $m_{\rm\chi}>m_{\rm h}/2$, production is dominated by 2-to-2 scatterings, which changes the slope of the $\Lambda$ vs. $m_{\rm\chi}$ relation (dashed line). Due to the complexity of the interactions, an analytical estimation of the slope in this regime is not possible. Note that when $w_{\rm\phi}>3/7$, all DM masses are produced before EWSB irrespective of $T_{\rm re}$, and its behavior is the same as the pre-EWSB case, therefore we do not show the parameter space for $w_{\rm\phi}=0.5$.
@@ III.2.2 Freeze-out
We have illustrated the behavior of the DM parameter space $(\Lambda,m_{\rm\chi})$ in Fig. 8 for the freeze-out mechanism in the post-EWSB scenario. This parameter space is obtained by numerically solving Eq. 45 together with Eq. 5. We find that any DM mass $m_{\rm\chi}\gtrsim\mathcal{O}(4)$ TeV freezes out before EWSB; otherwise, freeze-out occurs after EWSB. Thus, the EWSB scenario can significantly influence the DM freeze-out, if the DM mass is below $\mathcal{O}(4)$ TeV, for any $w_{\rm\phi}$ and $T_{\rm re}$. The impact of reheating on freeze-out production is shown with red dashed lines. The solid red line corresponds to the standard freeze-out scenario, in which DM freezes out in the radiation-dominated epoch.
@@ IV DM parameter space and its constraints
In the previous section, we provided a detailed analysis of DM production and the behavior of its parameter space. In this section, we extend our discussion to explore the complete DM parameter space along with its relevant constraints. Figure. 9 illustrates the complete DM parameter space, where we consider six distinct reheating temperatures: $T_{\rm re}=(T_{\rm BBN},10,10^{5},10^{9},10^{12},10^{15})$ GeV, covering the entire range from the minimum to the maximum reheating temperatures. The colored shaded regions represent different constraints. The red-shaded region is excluded by the constraints on the reheating temperature $T_{\rm re}$. The light blue region represents the parameter space where the Higgs-Majorana interaction is Planck-suppressed ($\Lambda>M_{\rm pl}$), supported by the WIMPZILLA hypothesis Kolb and Long (2017). Since $T_{\rm max}$ is the maximum temperature during reheating, the effective theory of DM described in Eq. 21 is valid only if $\Lambda\geq T_{\rm max}$ Garcia et al. (2020); Barman et al. (2024a); the brown-shaded region violates this condition. This constraint rules out the entire freeze-out parameter space (see right plot), while in the freeze-in scenario, only certain regions of the DM parameter space are excluded. Furthermore, the unitarity constraint $\Lambda>m_{\chi}$ imposes an additional constraint, excluding the green-shaded region. The magenta-shaded region is discarded  [^5] For the freeze-out case, the DM parameter space is also affected by universal gravitational production, especially at very high masses Haque et al. (2023). However, since the entire freeze-out scenario is ruled out by the condition $\Lambda>T_{\rm max}$, we have not included it. due to the gravitational overproduction of DM Clery et al. (2022); Haque and Maity (2023).For gravitational production, we have considered contributions from both inflaton and thermal bath gravitational annihilations. The gravitational production from the thermal bath can significantly contribute to the present DM relic abundance if $T_{\rm re}\geq\mathcal{O}(10^{12})$ GeV Clery et al. (2022). Additionally, DM masses below $\mathcal{O}(10^{-5})$ GeV are excluded by Lyman-$\alpha$ Iršič et al. (2017); Ballesteros et al. (2021) constraints for freeze-in (see Appendix-C).The small dots represent the points where the freeze-in and freeze-out solutions meet, i.e., where there is a smooth transition from freeze-in to freeze-out or vice versa. Below this point, DM attains chemical equilibrium and is produced thermally via the freeze-out mechanism, whereas above it, chemical equilibrium is never established, and DM is generated non-thermally via the freeze-in mechanism. The corresponding mass at this point is the maximum DM mass capable of producing the correct relic abundance. For masses above this value, the DM becomes underabundant today.
@@
Figure 10: Comparison of the freeze-in DM parameter space $(\Lambda^{-1},m_{\rm\chi})$ with the sensitivity of current and future detection experiments. The three red lines correspond to different reheating temperatures: $T_{\rm re}=(T_{\rm BBN},0.5,10)$ GeV. Shaded regions indicate experimental sensitivities: XENONnT (green), LZ (purple), XLZD (orange), and LHC constraints (magenta), while the brown shaded area represents excluded parameter space ($\Lambda<T_{\rm max}$). The blue dashed line ($\nu$ fog”) marks the neutrino background limit for direct DM detection.”